# Intrinsic Heisenberg-type lower bounds on spacelike hypersurfaces in general relativity

### v2

@@ Abstract
We prove a coordinate- and foliation-independent Heisenberg-type lower bound for quantum states strictly localized in geodesic balls of radius $r$ on spacelike hypersurfaces of arbitrary spacetimes (with matter and a cosmological constant). The estimate depends only on the induced Riemannian geometry of the slice; it is independent of the lapse, shift, and extrinsic curvature, and controls the canonical momentum variance/uncertainty $\sigma_{p}$ by the first Dirichlet eigenvalue of the Laplace–Beltrami operator (Theorem). On weakly mean-convex balls we obtain the universal product inequality $\sigma_{p}r\geq\hbar/2$. Under the same assumption, a vector-field Barta-type argument improves this universal floor to the scale-invariant bound $\sigma_{p}r\geq\pi\hbar/2$, which provides a universal, foliation-independent floor. Any further sharpening of the constant requires eigenvalue‑comparison results or other curvature‑sensitive methods.
@@
Heisenberg-type inequality; quantum mechanics on curved spacetime; spacelike hypersurfaces; spectral geometry
PACS numbers:
03.65.Ta, 04.62.+v, 04.20.-q, 04.20.Cv, 02.30.Tb.
@@ 1 Introduction
On a Riemannian manifold $(\Sigma,h)$, strict localization of a quantum state to a bounded domain $\Omega\subset\Sigma$ with Dirichlet boundary conditions entails a non-vanishing spectral cost for the momentum observable: once $\psi|_{\partial\Omega}=0$, the $L^{2}$-norm of its gradient cannot be arbitrarily small. In spectral terms, the intrinsic uncertainty is governed by the first Dirichlet eigenvalue of the Laplace–Beltrami operator on $\Omega$ [1, Ch. 3]. This paper applies this principle to canonical data on Cauchy slices of curved spacetimes, making it directly usable from a physics perspective. On any spacelike hypersurface $(\Sigma,h)$ in a curved spacetime, we prove the following lower bound for states strictly localized to geodesic balls $B_{\Sigma}(p,r)$ [^1] Here $\lambda_{1}$ denotes the first Dirichlet eigenvalue of the Laplace–Beltrami operator $\Delta_{h}$ on $B_{\Sigma}(p,r)$.
$\sigma_{p}(\psi)\;\geq\;\hbar\,\sqrt{\lambda_{1}},$
where $\sigma_{p}(\psi)$ denotes the momentum standard deviation (Theorem 3.1). The estimate is coordinate- and foliation-invariant: it depends only on the induced Riemannian data of the slice and is manifestly insensitive to lapse, shift, and extrinsic curvature. In particular, the variance decomposition (15) and the kinetic-energy window
$\hbar\|\nabla^{h}|\psi|\|_{L^{2}}\;\leq\;\sigma_{p}(\psi)\;\leq\;\hbar\|\nabla^{h}\psi\|_{L^{2}},$
cf. (18), separate the geometric contribution carried by $|\psi|$ from the fluctuations of the phase gradient in an orthonormal frame. Operationally, this gives a clean, gauge-invariant bridge between spectral geometry on $(\Sigma,h)$ and momentum uncertainty defined by canonical commutators on the slice.
Beyond the main bound, we obtain a universal Heisenberg-type product inequality $\sigma_{p}\,r\geq\hbar/2$}, on weakly mean-convex geodesic balls of radius $r$; see Corollary 3.6. The argument uses only the boundary-distance Hardy inequality, valid under the distributional superharmonicity of the distance to $\partial B_{\Sigma}(p,r)$ [2, 3, 4]. The constant $1/2$ is optimal and the bound is never attained, so it serves as a robust, geometry-free floor whenever only coarse boundary information is available.
Under the same weak mean-convexity hypothesis, Corollary 3.7 sharpens this baseline by a factor of $\pi$ via a vector-field version of Barta’s method, yielding the scale-invariant estimate (23) [5]. This improvement depends only on the sign of the radial Laplacian and remains coordinate- and foliation-independent within the present framework; in particular, no symmetry, stationarity, or vacuum assumption is required on the chosen slice.
When curvature information is available, classical comparison techniques immediately transfer to quantitative control of $\lambda_{1}$ and thus of $\sigma_{p}$ : see, e.g., Cheng-type eigenvalue comparison theorems [6, 7], small-ball asymptotics [8], and explicit radial analyses on spherically symmetric manifolds [9]. In constant positive curvature, our universal product floor coexists with stronger model-specific bounds derived in [10].
@@
Organization. Section 2 introduces the minimal geometric setting and the precise Dirichlet localization; in particular, Assumption 2.1 records the standing conditions for $B_{\Sigma}(p,r)$. Section 3 establishes the intrinsic Heisenberg bound (Theorem 3.1). The subsections 3.1 and 3.2 derive the universal Hardy-Heisenberg baseline and the Barta-type improvement. Section 4 closes with a discussion of scope, extensions (including spinors/forms), and computational aspects.
@@ 2 Minimal geometric setting
Let $(\mathcal{M},g)$ be a time-oriented $C^{2}$ Lorentzian manifold solving the Einstein field equations (with arbitrary stress-energy tensor and possibly $\Lambda$). Let $\Sigma\subset\mathcal{M}$ be a $C^{2}$ spacelike hypersurface with induced Riemannian metric $h$ and volume element $\mathrm{d}\mu_{h}$. Fix $p\in\Sigma$ and $r>0$. Since $\Sigma$ is spacelike, it admits a unique future-directed unit normal field $n$ along $\Sigma$ with g(n,n)=1g(n,n)=-1. The induced metric is $h(X,Y):=g(X,Y)$ for $X,Y\in T\Sigma$, and the Levi–Civita connection $\nabla^{h}$ on $(\Sigma,h)$ is the tangential projection of the spacetime connection $\nabla$. A tubular neighborhood of $\Sigma$ is obtained by the normal exponential map $\exp^{g}$: for some $\varepsilon>0$, the map
$\Phi:(-\varepsilon,\varepsilon)\times\Sigma\to\mathcal{M},\qquad(t,x)\mapsto\exp^{g}_{x}\!\big(t\,n(x)\big)$
is a diffeomorphism onto its image; the level sets $\Sigma_{t}:=\Phi(\{t\}\times\Sigma)$ define a local foliation by spacelike slices with $\Sigma_{0}=\Sigma$. In these (Gaussian normal) coordinates, the metric takes the block form
$g=-\mathrm{d}t^{2}+h_{t},$
where $h_{t}$ is a $C^{1}$ family of Riemannian metrics on $\Sigma$ with $h_{0}=h$. More generally, for an arbitrary local time [^2] We use $\{t=c\}$ to denote the level set $t^{-1}(c)$ of a (local) time function $t$; thus $\Sigma=\{t=0\}$} is the chosen spacelike slice. function $t$ with $\Sigma=\{t=0\}$, the $(3+1)$-decomposition reads
$g=-N^{2}\mathrm{d}t^{2}+h_{ij}(t,x)\big(\mathrm{d}x^{i}+\beta^{i}\mathrm{d}t\big)\big(\mathrm{d}x^{j}+\beta^{j}\mathrm{d}t\big),$
with lapse $N>0$ and shift vector field $\beta$. The extrinsic curvature (second fundamental form) of $\Sigma$ is
$K(X,Y):=-\,g(\nabla_{X}n,Y)=-\tfrac{1}{2}\,(\mathcal{L}_{n}h)(X,Y),$
and, in Gaussian normal coordinates, $\partial_{t}h\!\mid_{t=0}=-2K$. None of the extrinsic data $(K,N,\beta)$ will be used in our estimates below; the bound depends only on the intrinsic geometry of $(\Sigma,h)$ and is therefore coordinate- and foliation-independent.
@@
All domains are taken within $(\Sigma,h)$. We write $B_{\Sigma}(p,r)$ for the $h$-geodesic ball of radius $r$ centered at $p$. Throughout we choose $r$ strictly below the injectivity radius at $p$ so that $B_{\Sigma}(p,r)$ is a bounded Lipschitz domain; in particular, the Dirichlet Laplacian on $B_{\Sigma}(p,r)$ has compact resolvent. These minimal hypotheses are recorded in Assumption 2.1 and will be in force for the main intrinsic estimate.
@@
  • (i)
    $r$ is strictly less than the injectivity radius at $p$ and chosen so that $B_{\Sigma}(p,r)$ is a bounded Lipschitz domain; in particular, the Dirichlet Laplacian has compact resolvent. [^3] The injectivity radius at $p\in(M,g)$, denoted $\operatorname{inj}(p)$, is the supremum of $r>0$ such that the exponential map $\exp_{p}\!:\,B_{T_{p}M}(0,r)\to M$ is a diffeomorphism onto its image; equivalently, every geodesic starting at $p$ is minimizing up to length $r$.
  • (ii)
@@
No further assumption (e.g., symmetry, stationarity, vacuum, or asymptotic structure) is required for the main intrinsic estimate. For the Hardy baseline in Corollary 3.6, we additionally assume that $\partial B_{\Sigma}(p,r)$ is weakly mean-convex (in the distributional sense), which is equivalent to superharmonicity of the boundary distance on $B_{\Sigma}(p,r)$ [2].
@@ 2.1 Hilbert space on geodesic balls
We work on an $n$-dimensional (or $n=3$) Riemannian manifold $(\Sigma,h)$ and restrict attention to the geodesic ball $B_{\Sigma}(p,r)\subset(\Sigma,h)$. Let $\{\partial_{i}\}$ be a local coordinate basis and let $\{X_{a}\}_{a=1}^{n}$ be an $h$-orthonormal frame. Indices $i,j,k,\dots$ refer to coordinates (raised/lowered with $h$), while $a,b,c,\dots$ refer to the orthonormal frame (raised/lowered with $\delta$). The frames are related by a vielbein $e_{a}{}^{\,i}$ and its inverse $e^{a}{}_{\,i}$:
@@
All integrals and $L^{2}$ norms below are taken over $B_{\Sigma}(p,r)$ with respect to $d\mu_{h}$.
For $v=\sum_{a=1}^{n}v^{a}X_{a}(x)\in T_{x}\Sigma$ we set
@@
and, for vector fields $V(x)$ on $B_{\Sigma}(p,r)$,
$\|V\|_{L^{2}}$ $:=$ $\Big(\int_{B_{\Sigma}(p,r)}\|V(x)\|_{h}^{2}\,d\mu_{h}\Big)^{1/2},$
@@
For normalized states $\psi\in L^{2}(B_{\Sigma}(p,r),d\mu_{h})$ with $\|\psi\|_{L^{2}}=1$, we write expectation values as $\langle A\rangle:=\langle\psi,A\psi\rangle_{L^{2}}$.

### v4

@@ Abstract
In quantum theory on curved backgrounds, Heisenberg’s uncertainty principle is usually discussed in terms of ensemble variances and flat-space commutators. Here we take a different, preparation-based viewpoint tailored to sharp position measurements on spacelike hypersurfaces in general relativity. A projective localization is modeled as a von Neumann–Lüders projection onto a geodesic ball $B_{\Sigma}(r)$ of radius $r$ on a Cauchy slice $(\Sigma,h)$, with the post-measurement state described by Dirichlet data. Using DeWitt-type momentum operators adapted to an orthonormal frame, we construct a geometric, coordinate-invariant momentum standard deviation $\sigma_{p}$ and show that strict confinement to $B_{\Sigma}(r)$ enforces an intrinsic kinetic-energy floor. The lower bound is set by the first Dirichlet eigenvalue $\lambda_{1}$ of the Laplace–Beltrami operator on the ball, $\sigma_{p}\geq\hbar\sqrt{\lambda_{1}}$, and is manifestly invariant under changes of coordinates and foliation. A variance decomposition separates the contribution of the modulus $|\psi|$ from phase-gradient fluctuations and clarifies how the spectral geometry of $(\Sigma,h)$ controls momentum uncertainty. Assuming only minimal geometric information, weak mean-convexity of the boundary yields a universal, scale-invariant Heisenberg-type product bound, $\sigma_{p}r\geq\pi\hbar/2$, depending only on the proper radius $r$.
@@
Heisenberg-type inequality; quantum mechanics on curved spacetime; spacelike hypersurfaces; spectral geometry; von Neumann-Lüders projection;
@@ 1 Introduction
Historically, Kennard [1] was the first to adopt the standard deviation as a quantitative measure of uncertainty in quantum mechanics. However, neither he nor Heisenberg [2, 3] provided an explicit justification for this choice from the standpoint of experimental physics. In the textbook formulation, Kennard’s choice [1] is effectively encoded in the ensemble-based variance relation
$\sigma_{p}\sigma_{x}\;\geq\;\frac{\hbar}{2},$
which is obtained by measuring position and momentum on two identically $\psi$-prepared subensembles.
By contrast, single-slit diffraction of particles has long been a paradigmatic illustration of Heisenberg’s uncertainty relations and their significance in quantum measurement [2, 3]. A beam of particles is sent through a slit of width $\Delta x$ in an otherwise opaque diaphragm; behind the slit, the transmitted wave is diffracted and acquires a spread in transverse momentum that grows as $\Delta x$ is reduced. This is the standard heuristic picture used to introduce Heisenberg’s principle [2, 3]. If one regards passage through the slit itself as a preparation procedure, then the relevant quantity is the momentum spread of the transmitted beam. A finite momentum standard deviation $\sigma_{p}<\infty$ is obtained only if the wave function of the transmitted state vanishes at the edges of the interval of width $\Delta x$; otherwise $\sigma_{p}$ is not even defined. Under this natural regularity condition the optimal sharp inequality reads [4, 5]
$\sigma_{p}\,\Delta x\;\geq\;\pi\hbar.$
Mathematically, this bound is nothing but the one-dimensional Wirtinger inequality ([6], Ch. 1.7), the prototype of the wider class of Poincaré inequalities. Physically, it embodies a different viewpoint from Kennard’s: the preparation is implemented by a concrete projection (the slit), and the uncertainty relation links a geometric parameter $\Delta x$ to the resulting momentum spread.
These two inequalities,
$\sigma_{p}\sigma_{x}\;\geq\;\frac{\hbar}{2}\qquad\text{and}\qquad\sigma_{p}\Delta x\;\geq\;\pi\hbar,$
are therefore operationally inequivalent. In the Kennard relation [1], one first chooses a state $\psi$ and then reads off $\sigma_{x}$ and $\sigma_{p}$ from separate ensembles prepared in that same state; the relation constrains the spreads of position and momentum for a fixed $\psi$. In the single-slit setting, by contrast, the preparation itself is given by a sharp spatial restriction: the incoming state $\psi$ is projected onto the interval of width $\Delta x$ defined by the slit, and the transmitted ensemble is described by the von Neumann–Lüders update of $\psi$ [7, 8]. The corresponding inequality directly relates the geometric preparation parameter $\Delta x$ of the projected state to its momentum standard deviation $\sigma_{p}$. Neither inequality can be derived from the other; they capture complementary aspects of the incompatibility of position and momentum.
If one asks how Heisenberg’s principle should be formulated in a relativistic setting, this preparation-based point of view becomes particularly important. In regions where gravity is strong, any attempt to localize a particle (or an event) within a very small spacetime region requires large momenta and therefore high energy. According to Einstein’s field equations, such energy densities curve the spacetime metric and may even lead to horizon formation. Much of the literature on this interface between quantum mechanics and gravity proceeds by generalizing Kennard’s ensemble-based relation: one modifies the canonical commutation relations or the functional form of the position-momentum uncertainty bound, leading to a family of so-called Generalized Uncertainty Principles (GUPs) [9, 10, 11, 12, 13]. Structurally, these proposals keep the Kennard picture of preparation and measurement but deform the algebra or the bound.
If one wants spacetime geometry and gravitational backreaction to enter already at the level of the preparation step, it is natural to turn the logic around. The single-slit experiment suggests that a sharp preparation is most faithfully modeled by a projection onto a spatial region, not by an abstract choice of wave function. In flat space this region is an interval or a slit; in general relativity, the appropriate analogue is a spatial domain on a Cauchy slice. The basic idea of this work is therefore to take von Neumann–Lüders projections onto higher-dimensional spatial regions as the starting point and to let the geometry of those regions control the resulting momentum uncertainty.
Concretely, we work on a Riemannian manifold $(\Sigma,h)$ representing a spacelike hypersurface of a Lorentzian spacetime. A sharp localization of a quantum state to a bounded domain $\Omega\subset\Sigma$ with Dirichlet boundary conditions ($\psi|_{\partial\Omega}=0$) plays the role of an idealized projective preparation: the probability to find the particle outside $\Omega$ vanishes. From the spectral point of view, such strict localization has an unavoidable cost. Once Dirichlet data are imposed, the $L^{2}$-norm of the gradient of $\psi$ cannot be made arbitrarily small: the intrinsic uncertainty is governed by the first Dirichlet eigenvalue of the Laplace–Beltrami operator on $\Omega$ [14, Ch. 3]. By the Rayleigh–Ritz characterization, the first Dirichlet eigenvalue $\lambda_{1}(\Omega;h)$ is the minimal value of the Dirichlet energy
$\int_{\Omega}\|\nabla_{h}\psi\|_{h}^{2}\,d\mu_{h}$
among all normalized wave functions $\psi\in H^{1}_{0}(\Omega)$. Thus strict localization to $\Omega$ enforces a nonzero kinetic-energy floor set by $\lambda_{1}$. Since the geometric momentum variance defined in Section 2.2 is directly related to this kinetic energy, $\lambda_{1}$ emerges as the intrinsic spectral invariant that fixes the minimal momentum uncertainty compatible with localization to $\Omega$.
To connect this principle to general relativity, we apply it to geodesic balls on Cauchy slices. Let $(\Sigma,h)$ be a spacelike hypersurface in a curved spacetime and $B_{\Sigma}(p,r)$ the $h$-geodesic ball of proper radius $r$ around a point $p\in\Sigma$. For states that are strictly localized to $B_{\Sigma}(p,r)$ with Dirichlet data, we prove the intrinsic lower bound
$\sigma_{p}\;\geq\;\hbar\sqrt{\lambda_{1}(B_{\Sigma};h)},$
where $\lambda_{1}(B_{\Sigma};h)$ is the first Dirichlet eigenvalue of the Laplace–Beltrami operator $\Delta_{h}$ on the ball, and $\sigma_{p}$ is the corresponding geometric momentum standard deviation (Theorem 3.1). The estimate is intrinsic: it depends only on the induced Riemannian metric $h$ on the slice and is manifestly insensitive to the lapse, the shift, and the extrinsic curvature. Passing to a different slice (at a different time or in a different foliation) generally changes the geometry of $B_{\Sigma}(p,r)$ and hence the numerical value of $\lambda_{1}$, but the geometric form of the bound is the same on each slice.
The variance decomposition derived in Section 2.2 makes this connection between geometry and uncertainty more transparent. Writing $\psi=|\psi|e^{i\varphi}$ one obtains a kinetic-energy window
$\hbar\|\nabla^{h}|\psi|\|_{L^{2}}\;\leq\;\sigma_{p}\;\leq\;\hbar\|\nabla^{h}\psi\|_{L^{2}},$
which separates the contribution coming from the modulus $|\psi|$ from the fluctuations of the phase gradient in an orthonormal frame. Operationally, this provides a clean bridge between spectral geometry on $(\Sigma,h)$ and the momentum uncertainty defined by the canonical commutator algebra on the slice. In contrast to GUP models that postulate modified commutation relations, our construction stays within standard quantum mechanics on a fixed curved background; all “new physics” is carried by the geometry of the localization region.
In many situations, however, the detailed interior geometry of $B_{\Sigma}(p,r)$ is not known, while some information about its boundary is available. This includes, for instance, numerical relativity setups where one can control the proper size and mean curvature of an outer boundary, or gravitational settings where a marginally outer trapped surface provides a distinguished two-sphere but the interior metric remains uncertain. Under a weak mean-convexity assumption on the boundary (made precise in Assumption 2.1), the distance to $\partial B_{\Sigma}(p,r)$ is distributionally superharmonic. In that case the sharp boundary-distance Hardy inequality [15, 16, 17] applies and yields a purely geometric lower bound on the first Dirichlet eigenvalue in terms of the radius $r$. Combined with our intrinsic estimate, this leads to a universal Heisenberg-type product inequality
$\sigma_{p}\,r\;\geq\;\frac{\hbar}{2}$
for all strictly localized states (Corollary 3.6). This bound is never attained, which makes it a robust, geometry-free floor on $\sigma_{p}$ that depends only on the proper radius $r$ and the boundary, but not on the detailed curvature in the interior.
The same weak mean-convexity hypothesis allows one to go further. Using a vector-field version of Barta’s method [18], adapted to the radial distance on $B_{\Sigma}(p,r)$, we construct a drift field that probes only the sign of the radial Laplacian. This yields a Barta-type lower bound on the first Dirichlet eigenvalue and hence a sharper, still scale-invariant product inequality
$\sigma_{p}\,r\;\geq\;\frac{\pi\hbar}{2}$
for all strictly localized states (Corollary 3.7). The improvement by a factor of $\pi$ again involves no further assumption on the slice: it is coordinate- and foliation-independent and does not require symmetry, stationarity, or vacuum conditions. In a suitable minimax sense in three dimensions, this constant is best possible under the sole hypothesis of weak mean-convexity; any further sharpening must draw on more detailed curvature information.
Whenever such curvature information is available, classical comparison techniques from spectral geometry can be imported directly to control $\lambda_{1}$ and thereby $\sigma_{p}$. Cheng-type eigenvalue comparison theorems [19, 20, 21, 22] relate the first Dirichlet eigenvalue on a geodesic ball in $(\Sigma,h)$ to that on model spaces with prescribed Ricci curvature; small-ball asymptotics [23] give precise expansions in the regime where the ball radius is much smaller than the curvature radius; and explicit radial analyses on spherically symmetric manifolds [24] provide detailed control in highly symmetric situations. For manifolds of constant curvature, the universal product bounds obtained here are complemented by sharper curvature-dependent estimates previously derived in [25].
@@
Organization. Section 2 introduces the minimal geometric setting and the notion of strict Dirichlet localization to a geodesic ball; in particular, Assumption 2.1 records the standing hypotheses on $B_{\Sigma}(p,r)$. Section 3 establishes the intrinsic Heisenberg-type lower bound in terms of the first Dirichlet eigenvalue (Theorem 3.1). Subsections 3.1 and 3.2 then derive the universal Hardy–Heisenberg baseline and the Barta-type improvement, respectively. Section 4 presents examples and applications, and Section 5 closes with a discussion of scope and possible extensions.
@@ 2 Minimal geometric setting
Let $(M,g)$ be a time-oriented $C^{2}$ Lorentzian manifold solving the Einstein field equations (with arbitrary stress-energy tensor and possibly $\Lambda$), and let $\Sigma\subset M$ be a $C^{2}$ spacelike hypersurface with induced Riemannian metric $h$ and volume element $d\mu_{h}$. We fix $p\in\Sigma$ and $r>0$. Since $\Sigma$ is spacelike, it carries a unique future-directed unit normal field $n$ along $\Sigma$ with g(n,n)=1g(n,n)=-1, and the Levi–Civita connection $\nabla^{h}$ on $(\Sigma,h)$ is the tangential projection of the spacetime connection $\nabla$. In what follows only the intrinsic Riemannian data $(\Sigma,h)$ will enter. None of the extrinsic quantities (lapse, shift, or extrinsic curvature) appear in the estimates, so all bounds below are coordinate- and foliation-independent.
@@
All domains are taken within $(\Sigma,h)$. We write $B_{\Sigma}(p,r)$ for the $h$-geodesic ball of radius $r$ centered at $p$. Recall that the injectivity radius of $(\Sigma,h)$ at $p$, denoted $\operatorname{inj}_{\Sigma}(p)$, is the supremum of $r>0$ such that the Riemannian exponential map
$\exp^{h}_{p}\colon B_{T_{p}\Sigma}(0,r)\to\Sigma$
is a diffeomorphism onto its image; equivalently, every $h$-geodesic starting at $p$ is minimizing up to length $r$. Throughout we choose $r$ strictly below $\operatorname{inj}_{\Sigma}(p)$ so that $B_{\Sigma}(p,r)$ is a bounded (Lipschitz) domain; in particular, the Dirichlet Laplacian on $B_{\Sigma}(p,r)$ has compact resolvent. These minimal hypotheses are recorded in the following Assumption 2.1 and will be in force for the main intrinsic estimate.
@@
  1. (i)
    $r$ is strictly less than the injectivity radius $\operatorname{inj}_{\Sigma}(p)$ and chosen so that $B_{\Sigma}(p,r)$ is a bounded Lipschitz domain; in particular, the Dirichlet Laplacian has compact resolvent.
  2. (ii)
@@
No further assumption (e.g., symmetry, stationarity, vacuum, or asymptotic structure) is required for the main intrinsic estimate. For the Hardy and Barta baselines in Corollary 3.6 and 3.7, we additionally assume that the boundary $\partial B_{\Sigma}(p,r)$ of the ball is weakly mean-convex. For $C^{2}$ boundaries this is known to be equivalent to the boundary-distance function $d(x)=\operatorname{dist}_{h}(x,\partial B_{\Sigma}(p,r))$ being distributionally superharmonic on $B_{\Sigma}(p,r)$ (i.e., $\Delta_{h}d\geq 0$); see Definition 3.5 below and [15, Thm. 1.2] together with [16].
Remark 2.2.
Strict localization and Dirichlet data. The Dirichlet condition in Assumption 2.1 should be viewed as an idealized notion of strict localization or von Neumann–Lüders preparation to $B_{\Sigma}(p,r)$. Physically, wave packets will typically have small “tails” that extend beyond any given geodesic ball. The present bounds are nevertheless relevant in that they describe the limiting behaviour of momentum uncertainty for families of states whose $L^{2}$-mass becomes more and more concentrated in $B_{\Sigma}(p,r)$. From a mathematical viewpoint, Dirichlet data are natural because they yield a purely bulk representation of the variance in terms of the Dirichlet form of $\Delta_{h}$ and a discrete Dirichlet spectrum; for Neumann or Robin boundary conditions, boundary terms survive and the clean spectral control used in Theorem 3.1 breaks down.
@@ 2.1 Hilbert space on geodesic balls
In the intrinsic, slice-based formulation adopted here, geodesic balls play a distinguished role as localization regions. On a spacelike hypersurface $(\Sigma,h)$, a geodesic ball $B_{\Sigma}(p,r)$ is defined purely in terms of the Riemannian distance induced by $h$; it does not depend on any particular coordinate chart or foliation, and it automatically adapts to the local curvature of the slice. This makes geodesic balls the most natural choice for implementing sharp, covariant position measurements in curved spacetime, in contrast to coordinate boxes or “cubes”, whose geometry is tied to a specific parametrization of $\Sigma$.
From the spectral point of view, the relevance of geodesic balls is even more pronounced. As shown later in Theorem 3.1, the intrinsic lower bound on the canonical momentum uncertainty of a state that is strictly localized in a bounded region $\Omega\subset\Sigma$ is governed by the first Dirichlet eigenvalue of the Laplace–Beltrami operator on $\Omega$. In many geometric settings, Faber–Krahn-type comparison results imply that, among all domains of a given volume, geodesic balls minimize this first Dirichlet eigenvalue. Equivalently, for fixed volume, geodesic balls yield the smallest possible spectral cost of strict localization and therefore the weakest (i.e. most forgiving) intrinsic lower bound on the momentum uncertainty. Any other region with the same volume has a larger first Dirichlet eigenvalue and thus enforces a stronger Heisenberg-type constraint.
For our purposes, this means that geodesic balls provide a canonical “worst-case” reference geometry: if one can establish universal lower bounds for the momentum uncertainty on $B_{\Sigma}(p,r)$, then the corresponding bounds for other shapes of the same volume are automatically no better and typically stricter. At the same time, the round geometry of geodesic balls keeps the subsequent analysis technically manageable and well aligned with classical tools from spectral geometry. In the remainder of this subsection we therefore formulate the Hilbert space setting directly on $L^{2}\bigl(B_{\Sigma}(p,r),d\mu_{h}\bigr)$, taking strict Dirichlet localization to the geodesic ball as our basic notion of position uncertainty.
We work on an $n$-dimensional (or n=3n{=}3) Riemannian manifold $(\Sigma,h)$ and restrict attention to the geodesic ball $B_{\Sigma}(p,r)\subset(\Sigma,h)$. Let $\{\partial_{i}\}$ be a local coordinate basis and let $\{X_{a}\}_{a=1}^{n}$ be an $h$-orthonormal frame. Indices $i,j,k,\dots$ refer to coordinates (raised/lowered with $h$), while $a,b,c,\dots$ refer to the orthonormal frame (raised/lowered with $\delta$). The frames are related by a vielbein $e_{a}{}^{\,i}$ and its inverse $e^{a}{}_{\,i}$:
@@
All integrals and $L^{2}$ norms below are taken over $B_{\Sigma}(p,r)$ with respect to $d\mu_{h}$. For $v{=}\sum_{a=1}^{n}v^{a}X_{a}(x)\in T_{x}\Sigma$ we set
@@
and, for vector fields $Y(x)$ on $B_{\Sigma}(p,r)$,
$\|Y\|_{L^{2}}$ $:=$ $\Big(\int_{B_{\Sigma}(p,r)}\|Y(x)\|_{h}^{2}\,d\mu_{h}\Big)^{1/2},$
@@
Born probability measure. The Hilbert space on the geodesic ball is $L^{2}(B_{\Sigma}(p,r),d\mu_{h})$, where $d\mu_{h}$ is the Riemannian volume element of the induced metric $h$. For a normalized state $\psi$, the density $|\psi|^{2}d\mu_{h}$ is therefore the natural Born probability measure for position on $B_{\Sigma}(p,r)$, and all expectations are taken with respect to this measure.